OpenStudy (anonymous):

need help!!!! solve the system by substitution { 2x-y+z=-4 { z=5 {-2x+3y-z=-10 A. (-8,7,5) B. (-8,-7,5) C. (8,-7,5) D. (-8,-7,-5)

ganeshie8 (ganeshie8):

start by substituting z=5 in first equation

OpenStudy (anonymous):

so it would be 2x-y+5=-4 and the other one would be -2x+3y-2=-10

ganeshie8 (ganeshie8):

Perfect ! next, write them in two rows and add them vertically

OpenStudy (anonymous):

*-5

OpenStudy (anonymous):

2x-y+5=-4 2x+3y-5=-10

OpenStudy (anonymous):

@ganeshie8 now what do you do ?

OpenStudy (anonymous):

put z in 1st and 3rd equation after that add 1st and 3rd equation which will give you y put y and z in 1st equation then get x... (:

OpenStudy (anonymous):

yeah i did that and i got 2x-y=-9 -2x+3y=-5 but i dont get how to solve that from there thats what im stuck on

ganeshie8 (ganeshie8):

write them two rows and add them vertically 2x-y = -9 -2x+3y = -5 ------------------

OpenStudy (anonymous):

okay does that mean 2x cancels out ?

ganeshie8 (ganeshie8):

yup

OpenStudy (anonymous):

yes dear

OpenStudy (anonymous):

and the equation would be 2y=-14 correct ?

OpenStudy (anonymous):

y=-7

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

Yes ! divide 2 both sides, and you will have y

OpenStudy (anonymous):

okay so i got that which equation am i suppose to plug those into ?

OpenStudy (anonymous):

any of 1st or 3rd

ganeshie8 (ganeshie8):

you have y & z values, so u can find the remaining x value by pluggin it in any equation

ganeshie8 (ganeshie8):

pick one of below 2x-y = -9 -2x+3y = -5 cuz you just need to plug in y value to solve x

OpenStudy (anonymous):

so the answer to the whole question would be B correct ? (-8,-7,5)

ganeshie8 (ganeshie8):

yes ! good job !!

OpenStudy (anonymous):

thank you for helping me :)

ganeshie8 (ganeshie8):

np :)